This FORS2 news page is separated into two parts. First a 'timeline of instrument status' is outllined, together with a table that lists important changes to the instrument status (this is formally complete from 13/09/20). Then, an 'announcements' section includes items regarding future changes to the instrument status. (This later section also includes instrument status changes prior to 13/09/20 that have not been incorporated into the table.)
Timeline of instrument status
The following table describes the events affecting the quality or the measurement of data obtained with FORS2. Its content can be downloaded as a Comma Separated Values file. It is complete as of September 13th, 2020.
The Start Date column indicates when the event took place. Some events cannot be defined by a single instant, but extend over sometimes several days, weeks or months. Operation with the instrument may continue while the event occurs, in which case the Ongoing flag is set. If the event had ended, the Ongoing ? flag is lowered and End Date is given. The Type column indicates the characteristic(s) of the event: Earthquake, Maintenance, Upgrade, Element failure, Intervention, Pipeline installation at QC, Change of nominal values or thresholds, Modification of configuration file of this database, or Other. The Instrument Modes column indicates the affected mode. The column KPI indicates which quality control parameters identified as Key Performance Indicators (i.e., parameters following the most critical quantities) are affected; the column non-KPI indicates which other critical Quality Control parameters are affected. The Calibration Raw Types column indicates the raw types of the affected calibrations. The Calibration Breakpoint flag indicates if calibrations obtained before the event cannot be used for science data obtained after the event or if calibration data obtained after the event cannot be used with science data obtained before the event. The Comment column provides more details. The History column allows one to see possible past information on the entry.
The Quality Control pages for FORS2 can be accessed through the QC overview page
- July 2021: from 05/07/21 all questions regarding SM or VM observing should be submitted via the new ESO Operations Helpdesk
- JANUARY 2021: from 13/09/20 events affecting the instrument status are reported in the above table and not here.
- March 2020: data obtained on the night of 12/02/2020 lack several keywords that are required for data reduction due to a software problem aflficting those data. This issue is currently under investigation and we hope that the keywords can be re-insrted into the headers asap. This page will be updated as soon as we have more information.
- January 2020: a) A new section has been added to the FORS2 User Manual (UM) that documents problems with flux calibration of FORS2 at wavelengths bluer than 5000A. There are two issues: 1) the varying flux of the blue flat-field lamp; 2) stray light from the red flat-field lamp. Both of these issues combine to result in errors of larger than 20% in flux calibration in the worst cases. These issues affect both FORS2 and FORS1 historical and new data. An investigation is ongoing. b) A note has been added to the FORS2 Template Manual (TM) that suggests that for spectral observations the user should use imaging filters in acqusition and through-slit templates that are consistent in wavelength with the subsequent grism for science observations. This is to avoid observations being flagged as laser sensitive if they are not.
- December 2019: the FORS2 User Manual (UM) and Template Manual (TM) have been updated (versions P105.1 and P105 respectively); to correct the units of offsets in science templates. ALL science template offsets are in arc seconds and NOT pixels (as previously stated).
- November 2019: the annual FORS2 maintenance took place, together with a cleaning of the LADC.
- September 2019: it has recently been noted that the scale of the blue-sensitive EEV detector (used in VM) is incorrect. This means that pre-images taken with the blue EEV CCD should NOT be used with FIMS to make MOS/MXU OBs - unless their astrometry is first fixed. This issue will soon be corrected.
- August 2019: the FORS2 ETC (indeed in all Paranal instrument ETCs) now has a 'Turbluence Category' in place of the seeing input in older ETC versions. In the case of FORS2 this simply separates the seeing conditions into percentages of the night when a limiting seeing is realised (for a 1hr OB) and therefore is very similar to the previous seeing input. (The reason for this change is so that other [Adaptive Optics] instrument ETCs can take into account the coherence time required for specific observations.) These changes are explained here: https://www.eso.org/sci/observing/phase2/ObsConditions.MUSE.html. The user is also reminded of the change below w.r.t. PWV constraints. PWV can also now be set in the FORS2 ETC. By default this is set to the maximum value of 30mmm for FORS2, however users can set this to lower values if it is required by their science (for the vast majority of science this is not expected to be the case).
- May 2019: a) P104.1 version of the User Manual released with minor changes. b) all grisms are now permitted for use in Service Mode (SM). c) at phase2 users can now set the Precipitable Water Vapour (PWV) constraint for FORS2 OBs. By default this is set to the maximum value of 30mm for all FORS2 OBs. However, users can set this to lower values if it is required by their science.
- January 2019: ghost reflections have been found in imaging data obtained using the H_Alpha+83 filter. This also appears to affect historical data obtained with FORS2. An investigation is ongoing to characterise these features.
- November 2018: the annual FORS2 maintenance took place, together with a cleaning of the LADC.
- August 2018: new versions (P103) of the User Manual and Template Manual have been released. Specifically these try to make clearer how offsets are dealt with in both blind offsets in acquisition templates and offsets during science templates.
- July 2018: p2pp is replaced by the web-based p2 tool for all SM and VM observations, see: https://www.eso.org/sci/observing/phase2/p2intro.html. The P102.1 version of the User Manual (UM) no longer refers to p2pp but p2. In addition, a change has been made in the UM with respect to the on-sky region that is not observable when using the 600RI grism: from 21 to 35 arc seconds. The FIMS and Template UMs have also been updated: HIT-mode references have been removed, while p2pp references have been changed to p2.
- February 2018: the high time resolution modes (imaging, HIT-I; spectroscopy, HIT-S; and multi-object spectroscopy, HIT-MS) are no longer offered at FORS2 from period 102 onwards. All references to these modes have been removed from the public webpages: for use of these modes in P101 users should refer to the P101 User Manual and/or contact the FORS2 operations team directly (email@example.com, or firstname.lastname@example.org). All discussion of the HIT modes, their use and their characteristics have been removed from the P102 User Manual. For reference, the P101 User Manual - the last manual to include discussion of the HIT modes - will be kept on the documentation page.
- December 2017: a) filters FILT_388_4+118; FILT_621_5+87; and FILT753_8+89 all added to FORS2 ETC, b) filter characteristics (for same filters in a)) updated in User Manual; c) FIMS is updated for P101 Phase 2 to work on Mac machines. In addition to FIMS now being available for Mac OS (X) as well as Linux, it is now available via an official ESO MacPort repository for Mac OS (X) and official ESO RPM repositories for Fedora 20-25 and Scientific Linux 7. Fedora 26+, CentOS 7, RedHat Enterprise Linux 7 can also use these repositories and an Ubuntu repository is being trialed. A link to these repositories will be available soon after webletters are sent to users at the end of December 2017.
- October 2017: the exposure time calculator (ETC) for FORS2 spectral observations (https://www.eso.org/observing/etc/bin/gen/form?INS.NAME=FORS+INS.MODE=spectro) now displays a plot of the "Standard deviation of Stokes parameters (polarimetry)" in place of "Signal-to-noise" in the case of spectro-polarimetry (PMOS) observations. The standard deviation of the Stokes parameter is a more useful figure for PIs to estimate the exposure times if they know the degree (and nature) of the polarization of their target, and therefore after some internal investigation the ETC was changed in this direction. We also note that the use of the polarimetric mode (i.e. the Wollastom prism) now carries a 25% flux penalty in place of the previously used 50%. This was revised after the internal investigation.
- August 2017: In the CfP period 101, two additional narrow-band filters will be offered in VM at FORS2. These are 'FILT_753_8+89' (central wavelength of 7528A, FWHM ~80A), and 'FILT_621_5+87' (central wavelength of 6195A, FWHM ~60A). A full characterisation of these filters is ongoing, and potential users are encouraged to contact the FORS2 instrument team (email@example.com or firstname.lastname@example.org) in advance of proposal submission.
- August 2017: ESO is considering to decommission the high time resolution modes (imaging, HIT-I; spectroscopy, HIT-S; and multi-object spectroscopy, HIT-MS) currently offered at FORS2, due to their low demand and high overhead for the observatory. Potential future users of these modes are encouraged to contact the FORS2 instrument team (email@example.com, or firstname.lastname@example.org) to evaluate the scientific merit of the continued use of FORS2 HIT modes.
- August 2017: FORS2 User Manual updated for P101.
- 25 Dec 2016: FORS2 user manual updated.
- 1 Dec 2016: The FORS2 calibration plan has changed. The validity period for imaging sky flatfields has been increased to 14 days. If sky flats are not taken within this validity period then service mode OBs will typically be resheduled. The change in the calibration plan was the result of a study of the stability of imaging flatfields over the period of a month. Note that in normal operations then the current frequency of sky flats will likely be maintained.
- 14 Nov 2016: The use of the blue-optimized FORS1 detector E2V cannot be requested in Service Mode but may be requested in Visitor Mode. As of P98, observations requiring good blue sensitivity may select CCD=ANY in the acquisition template to allow execution with the E2V detector when it becomes available just a day or two before a scheduled visitor mode run using the same detector. Users selecting CCD=ANY should mention this in the Readme file.
- 12 September 2016: The primary mirror (M1) on Antu has been recoated. Preliminary analysis indicates that the R-band zero point has improved by approximately 0.16 magnitudes. See: http://www.eso.org/observing/dfo/quality/FORS2/reports/HEALTH/trend_report_ZEROPOINTS_R_HC.html
- 12 August 2015: Twilight flats obtained after the LADC prism exchange have now been analysed, confirming that the structure clearly visible in the old data is gone. The new data show a gradient across the field but no small-scale structure. Read the full report here.
- March 2015: A report on a project to improve the quality of the reduced FORS2 spectra by Sabine Moehler et al. is available in the ESO Messenger.
- 12 February 2015: A report presenting the prism exchange of the FORS2 LADC and its influence on exoplanet transits will soon appear in The Messenger. It is available from http://arxiv.org/abs/1502.03172 .
- 15 November 2014: The FORS2 LADC, which was suffering from a poor coating, was replaced by the one of FORS1 whose coating had been removed. This led to a gain in throughput of 0.1 ( in the blue), 0.07 (in V), 0.05 (in I), and 0.04 (in I) magnitudes, due to a reduction of the scattering.
- 24 August 2014: The FORS keyword HIERARCH ESO INS SLIT WIDTH (slit width in mm) was found wrong in the LSS, HIT and FREE mode, being a factor two too small. This affected all observations done with FORS1 and FORS2 till now. It has been corrected and all files in the archive corrected. The slit width in arcsec is correct.
- 24 June 2014: Since the last CCD exchange on 5 June and 22 June, FORS2 had a rotation misalignment of 0.024 radians. This could affect some runs, so Pis of runs executed during this period should take particular care.
- 3 June 2014: New tests confirm that cross-talk from linear to circular polarisation is present (confirming Bagnulo et al. 2009). Note that it appears only when we want to measure a weak field of circular polarisation in a source that is strongly linearly polarised. It is stronger in the blue (~10%, i.e., a signal of linear polarisation of the order of 10% is seen as a signal of circular polarisation of 1%) than in the red (in the R filter seems about 5%), but is not stable in time. Origin may be the LADC or the SR-collimator. Further tests are ongoing, but users should be cautious in interpreting their FORS2 polarisation data.
- 6 February 2014: The blue calibration lamps have been behaving erratically between 5 and 29 January, leading to underexposed spectroscopic flat fields in the blue (and thereby affecting grisms 300V, 600B and 1200B).
We have retaken when possible new calibrations, so users should check if their data are well calibrated.
- 06 November 2013: The coating of the LADC is severely worn off since 2 years. Programmes that require extremely precise photometry or spectrophotometry over several hours (e.g. transiting exoplanets) should thus not be attempted for now.
- 15 October 2013: Over the last year, we have had several instances where the retarder angle will go to an INDEF position - this will be in the header files (or sometimes nothing is written in the file) - and there is no way to know what was its actual position, although it seems to be mainly a software bug. This affects mostly previous visitor runs, and very seldomly, service mode runs.
- 01 October 2013: calibration OCA rules are now available for historical data, both for FORS1 and FORS2, back to March 2002. For these data, calselector will provide the necessary calibrations.
- 15 February 2013: It was discovered that due to a bug in the associated software, MOS calibrations taken since July 13, 2011 were systematically taken without slit #19. In other words, there has been no FLAT or WAVE taken for this slit since the above mentioned date.
- August 2012: from 10 to 15 of August the instrument was operated with the E2V (blue) mosaic, due to a failure in the control electronics of the MIT (red) mosaic.
- January 2010: the paper by Moehler et al.: Correction of Field Rotator-Induced Flat-Field Systematics has been published.
- April 2009: FORS1 decommissioned
- September 26, 2008: P83 proposers wishing to use grisms not covered bythe Standard Instrument Configuration should be aware that suchrequests will not be considered for Service Mode. For Visitor Modeproposals, users should propose for and base their time request on agrism covered by the Standard Instrument Configuration (found inSection 2.2 of the manual). In addition to this, they should include adescription of the gains that could be made from using a grismpreviously available on FORS (eg. 600V, 1200g). Such requests will beconsidered on a case-by-case basis during the technical evaluation.
- September 1, 2008: There was a error in Table 2.1 ofthe manual showing the "Standard Instrument Configuration". This has now been fixed and the information can also be found from the following link.
- August 29, 2008: Important information from the P83 Call for Proposals - "Firstly, the polarimetric capabilities previously offered on FORS1 will be transferred to FORS2. They will be available on FORS2 for observations during P83. Secondly, the blue-optimized E2V CCD that is currently on FORS1 will be available for Visitor Mode observations with FORS2 during the second half of P83. Finally, the standard instrument configuration for Service Mode observations with FORS2 has changed to incorporate a number of the high-throughput filters and VPH grisms previously used on FORS1". We will update this information as it becomes available during the semester. Users should consult the P83 FORS user manual, available here.
- March 1, 2008: Important extract from the P82 Call for Proposals - "Thepolarimetric capabilities of FORS1 will only be offered during thefirst half of P82. After this time, FORS1 will be offered in itsstandard configuration with the exception of the IPOL and PMOS modes.Due to the arrival of X-Shooter at the Cassegrain focus of UT2, it isexpected that FORS1 operations will end during the second half of P82."
- February 27, 2008: Updated FORS1 standardinstrument configuration - The R_BESS filter is moved to from Wheel 1to Wheel 3 and the I_BESS filter from Wheel 3 to Wheel 1. This meansR_BESS will be available for IPOL observations, but I_BESS will not.
- July 16, 2007: Links to documents relating to the "FORS absolute Photometry Project" added to web page here.
- May 21, 2007: FORS spectroscopic pipeline released. See the DFS pages for more details here.
- April 12, 2007: These webpages are updated to take into account the new blue FORS1 detector.
- April 1 to April 6, 2007: The second commissioning of the new bluesensitive E2V CCD on FORS1 is completed with success. The new cameraand new high throughput broad band filters are then already offered inP79.
- March 1, 2007: The new mosaic CCD will be offeredon FORS1 for the beginning of P80 at the latest, so all P80 proposersshould read the information at Blue upgrade to find details of the enhanced capabilities of the instrument. Inaddition the standard configuration of FORS1 will change from P80 toincorporate the new filters.
- January 30 to February 4, 2007: First commissioning of the new blue sensitive E2V CCD on FORS1
- November 28, 2006 to December 9, 2006: FORS1 annual maintenance (LADC cleaning, new calibration screen).
- November 28, 2006 to December 6, 2006: UT2 M1+M3 mirrors re-coating.
- April 1, 2006 : A new HIT mode, the 'multiple-shift' mode, is available on FORS2 for fast simultaneous spectroscopy of science and reference targets.
- April 1, 2006 : A new high through-put Volume Phased Holographic grating is offered for FORS1. The new 1200B grism covers the range 373-497nm with a dispersion of 0.61Å/pixel and replaces the 1200g grism in the FORS1 standard grism set.
- August 30, 2005 : As of P77, the FORS2 Echelle is decommissioned and replaced by the new and more efficient 1200B grism on FORS1.
- February 25, 2005 : These new FORS webpages are released
- December 13, 2004 : New Rapid Response Mode (RRM) running for both FORS1+2
- June 15, 2004 : New holographic grism GRIS_1200g installed on FORS1
- June 5, 2004 : FORS1 moved to Kueyen (UT2) and FORS2 to Antu (UT1)
- February 27, 2004: Important update on the FORS1 instrumental polarization